A crazed subway sandwich shop employee hacks an 8 foot submarine sandwich into four unequal pieces with a machete. The first piece is one half as long as the second piece, the third piece is 12 inches longer than the second, and the fourth piece is the same length as the second. How long is the longest piece of sandwich?
Answer by stanbon(75887)

Let length of 1st piece be "x"
Then length of 2nd is "2x"
3rd piece is "2x+12"
4th piece is same as 2nd or "2x"
These are all in inches so change the 8 ft. to 96 inches.
EQUATION:
length + length + length + length=96 inches
x+2x+2x+12+2x=96
7x+12=96
7x=84
x=12 inches (length of 1st piece)
longest piece is 2x+12=2(12)+12=36 inches
Cheers,
Stan H.